OpenStudy (anonymous):

this have to be true or false. 7<9

OpenStudy (anonymous):

True

OpenStudy (anonymous):

5>5

OpenStudy (anonymous):

true , and false its = to!

OpenStudy (anonymous):

4<3

OpenStudy (anonymous):

false 4 is greater than 3 and it says its less than so false

OpenStudy (anonymous):

9>7

OpenStudy (anonymous):

true this is easy you can do it!

OpenStudy (anonymous):

if 8 cases of merchandise cost $60, what would a dozen cases cost

OpenStudy (anonymous):

7.5

OpenStudy (anonymous):

So 7.50$

OpenStudy (anonymous):

8/n=1/2

OpenStudy (anonymous):

$90

OpenStudy (anonymous):

these are proportions

OpenStudy (anonymous):

and 8/n=1/2 so n = 4

OpenStudy (anonymous):

i got 16

OpenStudy (anonymous):

well you multiply 8 by both sides and 1/2x8=4

OpenStudy (anonymous):

so n is by itself